# Transientangelo: Few-Viewpoint Surface Reconstruction Using Single-Photon Lidar

### v1

@@ Abstract
We consider the problem of few-viewpoint 3D surface reconstruction using raw measurements from a lidar system. Lidar captures 3D scene geometry by emitting pulses of light to a target and recording the speed-of-light time delay of the reflected light. However, conventional lidar systems do not output the raw, captured waveforms of backscattered light; instead, they pre-process these data into a 3D point cloud. Since this procedure typically does not accurately model the noise statistics of the system, exploit spatial priors, or incorporate information about downstream tasks, it ultimately discards useful information that is encoded in raw measurements of backscattered light. Here, we propose to leverage raw measurements captured with a single-photon lidar system from multiple viewpoints to optimize a neural surface representation of a scene. The measurements consist of time-resolved photon count histograms, or transients, which capture information about backscattered light at picosecond time scales. Additionally, we develop new regularization strategies that improve robustness to photon noise, enabling accurate surface reconstruction with as few as 10 photons per pixel. Our method outperforms other techniques for few-viewpoint 3D reconstruction based on depth maps, point clouds, or conventional lidar as demonstrated in simulation and with captured data.
Figure 1: Transientangelo takes as input raw lidar scans from sparse viewpoints. These scans are used to optimize an SDF-based scene representation, which is further regularized to constrain the geometry from both captured and unseen viewpoints. The method recovers higher-fidelity surfaces than previous methods in the sparse-view and low-photon regime (i.e., from tens to hundreds of measured photons per pixel). The above scene was trained using simulated single-photon lidar data from five viewpoints, with an average of 150 photons per pixel over the occupied regions of the scene.
@@ 1 Introduction
Previous methods for multi-viewpoint 3D reconstruction—whether based on conventional images or lidar—have several drawbacks. First, image-based methods are usually trained with tens or hundreds of images, making acquisition difficult [35, 32, 51, 24]. While several recent works attempt to tackle the few-view reconstruction problem [36, 29, 50, 63], they typically fail for the case where the baseline between input images becomes large and the problem becomes increasingly ill-posed. Recent methods that leverage large-scale diffusion models [55, 10] alleviate this limitation to some extent, but are also prone to hallucinating scene geometry. Other techniques improve few-view 3D reconstruction by introducing additional geometric constraints from point clouds [45] or monocular depth estimators [63]. In this vein, our work is closest to TransientNeRF [29], which shows that using raw lidar data instead of depth maps or point clouds leads to significant improvements in reconstructed scene appearance in the few-view setting. However, TransientNeRF is based on a volumetric representation and is designed for appearance modeling rather than surface reconstruction. Further, TransientNeRF uses long acquisition times ($\approx$20 minutes per view) to capture per-pixel photon count histograms with several thousand photon arrivals each.
@@
Overall, we improve few-viewpoint 3D surface reconstruction from raw lidar measurements compared to previous work (see Figure 1), while also demonstrating significant improvements in reconstructed geometry compared to image-based or depth-based approaches. Our approach also improves robustness to lidar measurements captured with relatively few photon counts (e.g., 10–300), which is especially important when operating at fast acquisition speeds or when imaging targets at long range. We make the following technical contributions:
@@
  • We propose regularization techniques, that improve the reconstructed geometry and provide improved robustness in the low-flux regime where only tens to hundreds of photons are collected per pixel.
@@
2 Related Works
Figure 2: Transientangelo training procedure. For a pixel $\mathbf{p}$, we cast out a ray $\mathbf{r}(t)=\mathbf{o}+tc\bm{\omega}(\mathbf{p})$. Captured viewpoint: 3D ray coordinates are passed through the neural surface representation $\mathcal{F}$ to retrieve a radiance and an SDF value, which gets converted to density $\sigma$ (see Equation  3). These values are then binned into a transient, which after a convolution with the laser pulse gives the final rendered transient $\bm{\tau}_{f}[i,j,k]$. The network is supervised with an L1 loss between the rendered and captured transient and an L1 loss between the integrals of the rendered and captured transients (reflectivity loss). Unseen viewpoint: 3D ray coordinates are passed through the neural representation to retrieve rendering weights, which are used to calculate the variance of the weights around the depth $d$. The network is trained to minimize this variance, resulting in thinner surfaces and removing spurious zero-level sets.
@@ 2 Related Works
While many previous techniques on 3D imaging with SPADs use high-powered lasers or long acquisition times to capture hundreds or thousands of photons [29, 25, 26, 39], our work is more closely related to single-photon imaging techniques that operate in the photon-starved regime. Such methods exploit spatial or temporal correlations in the photon arrivals to recover depth or reflectivity from only a few captured photons at each pixel [21, 42]. Our approach exploits cross-viewpoint information and reflectivity-based regularization for scene reconstruction from as few as 10 collected photons per pixel.
@@
Neural surface representations typically use either occupancy [38, 37, 40] or a signed distance function (SDF) [51, 52, 24, 60, 61] to model scene geometry. The former uses a neural network to classify points in 3D space that fall inside an object [38]. The latter learns a network that predicts the SDF of a 3D point (i.e., its signed distance to the closest surface). The surface is extracted from the SDF by finding the zero-level set. Our work, inspired by Neuralangelo [24], builds on a hashgrid-based SDF representation  [35]. However, different from previous work, we use transient measurements as supervision.
@@
Prior work has addressed few-viewpoint reconstruction using geometric supervision (e.g., based on depth maps, point clouds, or transients) [29, 50, 63, 4, 46], or by introducing other strategies to regularize reconstruction from multi-view images [36, 63, 17, 20, 57, 62, 54, 59]. Our work is closest to TransientNeRF [29] and RegNeRF [36]. In contrast to TransientNeRF—which reconstructs a volumetric scene model from few-viewpoint transients—we focus on surface reconstruction using an SDF-based representation and improve robustness to noisy (low photon) measurements. Similar to RegNeRF, we regularize views rendered from novel viewpoints; however, we use a different depth-based regularization strategy and achieve improved performance in the few-view setting by exploiting transient data.
@@ 3.1 Measurement Model
Here, $\bm{\lambda}$ is the expected number of incident photons and depends on the reflectivity of the point imaged by pixel $(i,j)$. The value $N$ denotes the number of emitted laser pulses per pixel, $\eta$ is the detection efficiency of the sensor, and $B$ is the number of background detections, which dependens on the average ambient photon rate at pixel coordinates $(i,j)$, $A[i,j]$. Finally, the value $D$ is the dark count—i.e., the number of false detections produced by the sensor.
@@ 3.3 Optimization
Objective function.
As our main training objective, we supervise the rendered transient using the measurements $\mathbf{\tilde{\bm{\tau}}}$:
@@
To sample unseen viewpoints we follow RegNeRF [36] and sample camera origins on a sphere around the scene estimated from the training views (see supplement for details). We then regularize the density from these unseen views using RawNeRF’s weight-variance regularizer [31]. However, we find that replacing their rendered depth estimate with the argmax depth [29] helps recover more accurate surfaces, as shown in the supplement.
More specifically let $t_{i}$ be the samples along the ray used in the quadrature approximation of Eq. 4. Then the values $w_{i}=\sigma(t_{i})T(t_{i})$ represent alpha compositing weights applied to the radiance values along a ray. We use $w_{i}$ to compute the weight-variance regularizer, which we apply to rays $\mathcal{R}_{\text{train}}^{\prime}$ sampled from the set of unseen views:
@@
Here, $d=\operatorname*{arg\,max}_{t^{*}}\;T(t^{*})\sigma(t^{*})$ represents the depth, or distance along the ray with the maximum probability of ray termination [29].
@@
To improve performance in the low photon count regime, we introduce a reflectivity loss, which provides additional supervision using the integrated transients. Since the integrated transients have a higher signal-to-noise ratio than the individual photon count histogram bins, this provides a complementary source of supervision.
We implement this loss as
$\mathcal{L}_{\text{ref}}=\frac{1}{|\mathcal{R}_{\text{train}}|}\sum_{\mathcal{R_{\text{train}}}}\|\sum_{n}\bm{\tau}[i,j,n]-\sum_{n}\tilde{\bm{\tau}}[i,j,n]\|_{1}.$ (7)
In summary, our training loss can be expressed as the following:
@@
$\quad+\lambda_{\text{weight\_var}}\mathcal{L}_{\text{weight\_var}}+\lambda_{\text{sparse}}\mathcal{L}_{\text{sparse}},$ (8)
where
$\mathcal{L}_{\text{sc}}=\sum\limits_{\begin{subarray}{c}i,j,n\\ \bm{\widetilde{\tau}}[i,j,n]<B\end{subarray}}\int_{\mathcal{P}_{i,j}}\int_{\mathcal{T}_{n}}T(t)\sigma(\mathbf{r})\,\mathrm{d}t\,\mathrm{d}\mathbf{p},$ (9)
is the space carving loss [29], which penalizes densities at locations where the transient is below a background level $B$.
@@
is the Eikonal regularization [12], which encourages the SDF to have gradient magnitude equal to one, and
@@
is the sparsity loss, which helps to prevent floating artifacts, and $\alpha$ is a scalar hyperparameter.
@@ 3.4 Implementation Details
We initialize the SDF to approximate a sphere [2]. We follow Neuralangelo’s coarse-to-fine optimization strategy by initializing the step-size in the gradient calculation $\epsilon$ as the coarsest hash-grid voxel size and exponentially decreasing it during training. For the hash-grid, we start at 4 hash resolutions and unmask two new hash resolutions every 5000 steps. We optimize the network using the AdamW [27] optimizer, with $\beta_{1}=0.9$ and $\beta_{2}=0.99$ and employ a two-step learning rate scheduler by initializing the learning rate to $1\times 10^{-5}$ and linearly increasing it until it reaches $1\times 10^{-3}$ by the end of the 5000 steps. Then, the learning rate is exponentially decreased over the remaining steps, with a total of 250K steps. All experiments were run on a single Nvidia RTX A6000 GPU with 48GB of memory. We provide additional implementation details, including hyperparameter settings, in the supplement.
@@ 4 Multiview Transient Dataset
To test the robustness of different methods on low-SNR transient data we adapt the simulated and captured datasets introduced by TransientNeRF [29]. For each captured and simulated scene, we synthesize transients with an average of 300, 150, 50 and 10 photons per occupied (non-background) pixel. The paragraphs below outline how these datasets are created.
@@
To synthesize transients at different photon counts in simulation we match the flux from the time-resolved Mitsuba renderer [29] to the target number of photons per pixel. We then follow Malik et al. [29] and set the background flux to $0.1$ photons per $2850$ captured scene photons, which is set to match the statistics of the captured dataset. The flux is then Poisson sampled to simulate photon counts.
@@ 5 Results
Figure 3: Results on the simulated dataset using an average of 1500 photons per occupied (non-background) pixel. We show the recovered meshes from the baselines and the proposed method. For the transient-based methods, we also show rendered transients for the indicated pixels. Our method recovers smoother meshes with fewer missing parts. We also recover transients that better match the ground truth.
Table 1: Simulated results assessing image quality, depth accuracy and mesh quality. The left column indicates average photons per pixel.
@@
Table 2: Captured results assessing image quality, depth accuracy and mesh quality. The left column indicates average photons per pixel.
@@
We evaluate our method for transient and intensity image synthesis from novel viewpoints, depth rendering, and mesh recovery. We also benchmark the performance of our method’s robustness to low-SNR data. A more comprehensive set of results and ablation studies can be found in the supplement.
@@
To measure the mesh reconstruction quality we use Chamfer Distance. For image reconstruction evaluation we use PSNR and LPIPS [64]. We also evaluate the quality of the rendered transients using Transient IOU [29], defined as the intersection over the union of the rendered transient and the ground truth transient (see supplement for details).
@@ 5.1 Simulated Results
Figure 3 shows the recovered meshes and rendered transients for specific pixels from the proposed method compared to baselines. We show results for measurements with an average of 1500 photons per occupied pixel (results for other photon levels are included in the supplement). Due to a lack of any depth-based priors, Neuralangelo [24] and RegNeRF [36] both fail to recover accurate scene geometry. MonoSDF [63] recovers noisier surfaces, than the proposed approach, especially when trained with fewer input viewpoints. Compared to TransientNeRF [29], the proposed method recovers smoother surfaces with fewer missing parts and fewer floating artifacts, especially in the case of training with two or three input viewpoints. The results highlight the benefit of training a neural surface representation with transient supervision.
Figure 5: Novel view synthesis for varying photon levels. We show the rendered novel view on the cinema scene, trained on five viewpoints with an average of $50,300$ photons. We also show peak-time visualizations [28], which show the full transient in a single visualization. Hue encodes the time of peak intensity, brightness is modulated by the maximum intensity, and each band corresponds to an isochrone, or wavefront of equal path length. We show transient plots (right) for the pixel indicated on the ground-truth image.
Table 1 shows the quantitative performance of the proposed method against the baselines averaged across all scenes in the two, three, and five training viewpoint settings. We benchmark performance for varying measurement SNRs by training using transients with lower photon counts. To compute the metrics, we first scale the rendered transients so that the average photon count per occupied pixel (across all viewpoints) is the same across all photon levels.
The recovered meshes from the proposed method achieve significantly lower Chamfer Distance compared to the other methods and roughly 5×5$\times$5 × lower than the second-best method, TransientNeRF. The recovered depth maps are also more accurate than the baseline methods, and the image metrics are comparable. Finally, the transients rendered from our method are more similar to the ground truth based on the Transient IOU metric.
Table 3: Ablation study on the proposed method in the Lego scene. We present results without the sparsity loss (w/o $\mathcal{L}_{\text{sparse}}$), without the space carving loss (w/o $\mathcal{L}_{\text{sc}}$) and without the weight variance regularizer (w/o $\mathcal{L}_{\text{weight\_var}}$).

### v4

@@ Abstract
We consider the problem of few-viewpoint 3D surface reconstruction using raw measurements from a lidar system. Lidar captures 3D scene geometry by emitting pulses of light to a target and recording the speed-of-light time delay of the reflected light. However, conventional lidar systems do not output the raw, captured waveforms of backscattered light; instead, they preprocess these data into a 3D point cloud. Since this procedure typically does not accurately model the noise statistics of the system, exploit spatial priors, or incorporate information about downstream tasks, it ultimately discards useful information that is encoded in raw measurements of backscattered light. Here, we propose to leverage raw measurements captured with a single-photon lidar system from multiple viewpoints to optimize a neural surface representation of a scene. The measurements consist of time-resolved photon count histograms, or transients, which capture information about backscattered light at picosecond time scales. Additionally, we develop new regularization strategies that improve robustness to photon noise, enabling accurate surface reconstruction with as few as 10 photons per pixel. Our method outperforms other techniques for few-viewpoint 3D reconstruction based on depth maps, point clouds, or conventional lidar as demonstrated in simulation and with captured data.
Figure 1: Transientangelo takes as input raw lidar scans from sparse viewpoints. These scans are used to optimize a scene representation based on a signed distance function, which is further regularized to constrain the geometry from both captured and unseen viewpoints. The method recovers higher-fidelity surfaces than previous methods in the sparse-view and low-photon regime (i.e., from tens to hundreds of measured photons per pixel). The above scene was trained using simulated single-photon lidar data from five viewpoints, with an average of 150 photons per pixel over the occupied regions of the scene.
@@ 1 Introduction
Previous methods for multi-viewpoint 3D reconstruction—whether based on conventional images or lidar—have several drawbacks. First, image-based methods are usually trained with tens or hundreds of images, making acquisition difficult [36, 33, 52, 25]. While several recent works attempt to tackle the few-view reconstruction problem [37, 30, 51, 64], they typically fail for the case where the baseline between input images becomes large, and the problem becomes increasingly ill-posed. Recent methods that leverage large-scale diffusion models [56, 10] alleviate this limitation to some extent, but are also prone to hallucinating scene geometry. Other techniques improve few-view 3D reconstruction by introducing additional geometric constraints from point clouds [46] or monocular depth estimators [64]. In this vein, our work is closest to TransientNeRF [30], which shows that using raw lidar data instead of depth maps or point clouds leads to significant improvements in reconstructed scene appearance in the few-view setting. However, TransientNeRF is based on a volumetric representation and is designed for appearance modeling rather than surface reconstruction. Further, TransientNeRF uses long acquisition times ($\approx$20 minutes per view) to capture per-pixel photon count histograms with several thousand photon arrivals each.
@@
Overall, we improve few-viewpoint 3D surface reconstruction from raw lidar measurements compared to previous work (see Figure 1), while also demonstrating significant improvements in reconstructed geometry compared to image-based or depth-based approaches. Our approach also improves robustness to lidar measurements captured with relatively few photon counts (e.g., 10–300), which is especially important when operating at fast acquisition speeds or when imaging targets at long range. In summary, we make the following technical contributions.
@@
  • We propose regularization techniques that improve the reconstructed geometry and provide improved robustness in the low-flux regime where only tens to hundreds of photons are collected per pixel.
@@
2 Related Work
Figure 2: Transientangelo training procedure. For a pixel $\mathbf{p}$, we cast out a ray $\mathbf{r}(t)=\mathbf{o}+tc\bm{\omega}(\mathbf{p})$. Captured viewpoint: 3D ray coordinates are passed through the neural surface representation $\mathcal{F}$ to retrieve a radiance and an SDF value, which gets converted to density $\sigma$ (see Equation  3). These values are then binned into a transient, which, after a convolution with the laser pulse, gives the final rendered transient $\bm{\tau}_{f}[i,j,k]$. The network is supervised with an L1 loss between the rendered and captured transient and an L1 loss between the integrals of the rendered and captured transients (reflectivity loss). Unseen viewpoint: 3D ray coordinates are passed through the neural representation to retrieve rendering weights, which are used to calculate the variance of the weights around the depth $d$. The network is trained to minimize this variance, resulting in thinner surfaces and removing spurious zero-level sets.
@@ 2 Related Work
While many previous techniques on 3D imaging with SPADs use high-powered lasers or long acquisition times to capture hundreds or thousands of photons [30, 26, 27, 40], our work is more closely related to single-photon imaging techniques that operate in the photon-starved regime. Such methods exploit spatial or temporal correlations in the photon arrivals to recover depth or reflectivity from only a few captured photons at each pixel [22, 43]. Instead, our approach exploits cross-viewpoint information and reflectivity-based regularization for scene reconstruction from as few as 10 collected photons per pixel.
@@
Neural surface representations typically use either occupancy [39, 38, 41] or a signed distance function (SDF) [52, 53, 25, 61, 62] to model scene geometry. The former uses a neural network to classify points in 3D space that fall inside an object [39]. The latter learns a network that predicts the SDF of a 3D point (i.e., its signed distance to the closest surface). The surface is extracted from the SDF by finding the zero-level set. Our work, inspired by Neuralangelo [25], builds on a hashgrid-based SDF representation  [36]. However, unlike in previous work, we use transient measurements as supervision.
@@
Prior work has addressed few-viewpoint reconstruction using geometric supervision (e.g., based on depth maps, point clouds, or transients) [30, 51, 64, 4, 47], or by introducing other strategies to regularize reconstruction from multi-view images [37, 64, 17, 21, 58, 63, 55, 60]. Our work is closest to TransientNeRF [30] and RegNeRF [37]. In contrast to TransientNeRF—which reconstructs a volumetric scene model from few-viewpoint transients—we focus on surface reconstruction using an SDF-based representation and improve robustness to noisy (low photon) measurements. Similar to RegNeRF, we regularize views rendered from novel viewpoints; however, we use a different depth-based regularization strategy and exploit transient data to achieve improved performance in the few-view setting.
@@ 3.1 Measurement Model
Here, $\bm{\lambda}$ is the expected number of incident photons and depends on the reflectivity of the point imaged by pixel $(i,j)$. The value $N$ denotes the number of emitted laser pulses per pixel, $\eta$ is the detection efficiency of the sensor, and $B$ is the number of background detections, which depends on the average ambient photon rate at pixel coordinates $(i,j)$, $A[i,j]$. Finally, the value $D$ is the dark count—i.e., the number of false detections produced by the sensor.
@@ 3.3 Optimization
Data term.
We supervise the rendered transients to be consistent with the measurements $\mathbf{\tilde{\bm{\tau}}}$ by minimizing the following loss function.
@@
To sample unseen viewpoints, we follow RegNeRF [37] and sample camera origins on a sphere around the scene estimated from the training views (see supplement for details). We then regularize the density from these unseen views using RawNeRF’s weight-variance regularizer [32]. We find that replacing their rendered depth estimate with the argmax depth [30] recovers more accurate surfaces, as shown in the supplement.
More specifically, let $t_{i}$ be the samples along the ray used in the quadrature approximation of Eq. 4. Then, the values $w_{i}=\sigma(t_{i})T(t_{i})$ represent alpha compositing weights applied to the radiance values along a ray. We use $w_{i}$ to compute the weight-variance regularizer, which we apply to rays $\mathcal{R}_{\text{train}}^{\prime}$ sampled from the set of unseen views:
@@
The value $d=\operatorname*{arg\,max}_{t^{*}}\;T(t^{*})\sigma(t^{*})$ represents the depth, or the distance along the ray with the maximum probability of ray termination [30].
@@
To improve performance in the low photon count regime, we introduce a reflectivity loss, which provides additional supervision using the integrated transients. Since the integrated transients have a higher signal-to-noise ratio than the individual photon count histogram bins, this provides a complementary source of supervision. We implement this loss as
$\mathcal{L}_{\text{ref}}=\frac{1}{|\mathcal{R}_{\text{train}}|}\sum_{\mathcal{R_{\text{train}}}}\|\sum_{n}\bm{\tau}[i,j,n]-\sum_{n}\tilde{\bm{\tau}}[i,j,n]\|_{1},$ (7)
where the integrated transients are computed by summing over $n$.
In summary, our training loss can be expressed as
@@
$\quad+\lambda_{\text{weight\_var}}\mathcal{L}_{\text{weight\_var}}+\lambda_{\text{sparse}}\mathcal{L}_{\text{sparse}}.$ (8)
Here, the space carving loss, $\mathcal{L}_{\text{sc}}$, penalizes densities at locations where the transient is below a background level $B$:
$\mathcal{L}_{\text{sc}}=\sum\limits_{\begin{subarray}{c}i,j,n\\ \bm{\widetilde{\tau}}[i,j,n]<B\end{subarray}}\int_{\mathcal{P}_{i,j}}\int_{\mathcal{T}_{n}}T(t)\sigma(\mathbf{r})\,\mathrm{d}t\,\mathrm{d}\mathbf{p}.$ (9)
The Eikonal regularization [12] is given as
@@
and encourages the SDF to have a gradient magnitude equal to one. Finally, the sparsity loss is
@@
where $\alpha$ is a scalar hyperparameter. This loss helps to prevent floating artifacts in the reconstruction.
@@ 3.4 Implementation Details
We initialize the SDF to approximate a sphere [2]. We follow Neuralangelo’s coarse-to-fine optimization strategy by initializing the step-size in the gradient calculation $\epsilon$ as the coarsest hash-grid voxel size and exponentially decreasing it during training. For the hash-grid, we start at four hash resolutions and unmask two new hash resolutions every 5000 steps. We optimize the network using the AdamW [28] optimizer, with $\beta_{1}=0.9$ and $\beta_{2}=0.99$ and employ a two-step learning rate scheduler by initializing the learning rate to $1\times 10^{-5}$ and linearly increasing it until it reaches $1\times 10^{-3}$ by the end of the 5000 steps. Then, the learning rate is exponentially decreased over the remaining steps, with a total of 250K steps. All experiments were run on a single Nvidia RTX A6000 GPU with 48GB of memory. We provide additional implementation details, including hyperparameter settings, in the supplement.
@@ 4 Multiview Transient Dataset
To test the robustness of different methods on low-SNR transient data, we adapt the simulated and captured datasets introduced by TransientNeRF [30]. For each captured and simulated scene, we synthesize transients with an average of 300, 150, 50, and 10 photons per occupied (non-background) pixel. The paragraphs below outline how these datasets are created.
@@
To synthesize transients at different photon counts in simulation, we match the flux from the time-resolved Mitsuba renderer [30] to the target number of photons per pixel. We then follow Malik et al. [30] and set the background flux to $0.001$ photons per $2850$ captured scene photons, which is set to match the statistics of the captured dataset. The flux is then Poisson sampled to simulate photon counts.
@@ 5 Results
Figure 3: Results on the simulated dataset using an average of 6000 photons per occupied (non-background) pixel. We show the recovered meshes from the baselines and the proposed method. For the transient-based methods, we also show rendered transients for the indicated pixels. Our method recovers smoother meshes with fewer missing parts. We also recover transients that better match the ground truth.
Table 1: Simulated results assessing image quality, depth accuracy, and mesh quality. The left column indicates average photons per pixel.
@@
Table 2: Captured results assessing image quality, depth accuracy, and mesh quality. The left column indicates average photons per pixel.
@@
We evaluate our method for transient and intensity image synthesis from novel viewpoints, depth rendering, and mesh recovery. We also benchmark the performance of our method’s robustness to low-SNR data. Our supplementary contains a more detailed set of results and ablation studies.
@@
To measure the mesh reconstruction quality, we use Chamfer Distance. For image reconstruction evaluation, we use PSNR and LPIPS [65]. We also evaluate the quality of the rendered transients using Transient IOU [30], defined as the intersection over the union of the rendered transient and the ground truth transient (see supplement for details).
@@ 5.1 Simulated Results
Figure 3 shows the recovered meshes and rendered transients for specific pixels from the proposed method compared to baselines. We show results for measurements with an average of 6000 photons per occupied pixel (results for other photon levels are included in the supplement). Due to a lack of any depth-based priors, Neuralangelo [25] and RegNeRF [37] both fail to recover accurate scene geometry. MonoSDF [64] recovers noisier surfaces than the proposed approach, especially when trained with fewer input viewpoints. Compared to TransientNeRF [30], the proposed method recovers smoother surfaces with fewer missing parts and fewer floating artifacts, especially in the case of training with two or three input viewpoints. The results highlight the benefit of training a neural surface representation with transient supervision.
Figure 5: Novel view synthesis for varying photon levels. We show the rendered novel view on the cinema scene, trained on five viewpoints with an average of $50,300$ photons. We also show peak-time visualizations [29], which show the full transient in a single visualization. Hue encodes the time of peak intensity, brightness is modulated by the maximum intensity, and each band corresponds to an isochrone, or wavefront of equal path length. We show transient plots (right) for the pixel indicated on the ground-truth image (blue dot).
Table 1 shows the quantitative performance of the proposed method against the baselines averaged across all scenes in the two, three, and five training viewpoint settings. We benchmark performance for varying measurement SNRs by training using transients with lower photon counts. To compute the metrics, we first scale the rendered transients so that the average photon count per occupied pixel (across all viewpoints) is the same across all photon levels. The recovered meshes from the proposed method achieve significantly lower Chamfer Distance compared to the other methods and roughly 5×5$\times$5 × lower than the second-best method, TransientNeRF. The recovered depth maps are also more accurate than the baseline methods, and the image metrics are comparable. Finally, the transients rendered from our method are more similar to the ground truth based on the Transient IOU metric.
Table 3: Ablation study on the Lego scene with 6000ppp. We present results without the sparsity loss (w/o $\mathcal{L}_{\text{sparse}}$), without the space carving loss (w/o $\mathcal{L}_{\text{sc}}$), without the weight variance regularizer (w/o $\mathcal{L}_{\text{weight\_var}}$) and without the reflectivity loss (w/o $\mathcal{L}_{\text{ref}}$).
@@
*For this ablation we set the Eikonal loss to 00 and the weight variance loss to $10^{-5}$.